OpenStudy (anonymous):

Fill in the blank to make the expression a perfect square. y^2 - 8y + _

OpenStudy (anonymous):

+ 16

jimthompson5910 (jim_thompson5910):

take half of the y coefficient (-8) to get -8*(1/2) = -4 then square -4 to get (-4)^2 = (-4)*(-4) = 16

jimthompson5910 (jim_thompson5910):

so 16 goes in that blank